OpenStudy (anonymous):

4y + 3x = 5 for x How do I solve for the indicated variable?

jimthompson5910 (jim_thompson5910):

\[\large 4y + 3x = 5\] \[\large 3x = 5-4y\] \[\large x = \frac{5-4y}{3}\]